OpenStudy (anonymous):

Compare .1/3 4/6 Use <, >, or =.

OpenStudy (anonymous):

Lets simplify 4/6 = 2/3 . 1/3 < 2/3

OpenStudy (gorv):

4/6=2/3

OpenStudy (anonymous):

A class is selling magazines as a fundraiser. Of the 200 magazines sold, Ananda 1/8 sold of them. Gina sold 0.065 of the magazines. Ben sold 3/100 of the magazines, and Fatima sold 0.105 of the magazines. Who sold the most magazines?

OpenStudy (anonymous):

1/8 * 200 = 25 Magazines ; Ananda 0.065 * 200 = 13 Magaziens ; Gina 0.03 * 200 = 6 Magazines ; Ben 0.105 * 200 = 21 Magazines ; Fatima Ananda sold the most magazines